OpenStudy (anonymous):

a particle moves along the x-axis so that at time t >(or equal to) 0 its position is given by x(t)=2t^3 - 21t^2 + 72t-53. At what time t is the particle at rest?

OpenStudy (bahrom7893):

Particle is at rest when Velocity = 0

OpenStudy (bahrom7893):

V = X ' ( t ) = 6t^2 - 42t + 72 = 0

OpenStudy (bahrom7893):

V = t^2 - 7t + 12 = 0

OpenStudy (bahrom7893):

Solve for t: t1 = 3 t2 = 4 Particle is at rest at t = 3 and t = 4. Fan me if I helped, thanks!